OpenStudy (timaashorty):

I'm stuck with this one ;c. Please help. Suppose a parabola has vertex (5,-3) and also passes through the point (6,1). Write the equation of the parabola in vertex form. A. Y=(x-5)^2-3 B. Y=4(x-5)^2-3 C. Y=4(x+5)^2-3 D. Y=4(x-5)^2+3

OpenStudy (agent0smith):

Well first use the vertex... (5, -3) means y=a(x - 5) - 3 now use the point (6, 1) - plug in x=6 and y=1, and find a.

OpenStudy (timaashorty):

Now I know it's either A or B. I haven't learned how to find a, but i think a equals four?

OpenStudy (agent0smith):

y=a(x - 5) - 3 now use the point (6, 1) - plug in x=6 and y=1, and find a... 1=a(6 - 5) - 3 ^find a

OpenStudy (timaashorty):

a is 1, that means the correct answer is A.

OpenStudy (timaashorty):

cuz 6-5 is 1. then multiply a by one, you get a, or 1. Am i right?

OpenStudy (agent0smith):

no... you have to SOLVE for a. 1=a(6 - 5) - 3 6-5 is 1, so it becomes 1=a - 3

OpenStudy (timaashorty):

but that's not from my choices..

OpenStudy (agent0smith):

I know you can solve that for a... try it.

OpenStudy (timaashorty):

just tell me what's a-3. lol

OpenStudy (agent0smith):

haha i know you can do it ;) 1=a - 3 try adding 3 to both sides.

OpenStudy (timaashorty):

ohhhh ohh it's 4. !!

OpenStudy (timaashorty):

it's either 4, or 0. lol

OpenStudy (agent0smith):

Its four haha :) good job, i knew you could do it.

OpenStudy (timaashorty):

Yay! You'll be a great teacher, cuz you help figure it out before telling them the answer. Thank you very much agent Smith!

OpenStudy (timaashorty):

unless you're already a teacher, lol

OpenStudy (agent0smith):

Thanks... I am already a teacher :D I teach algebra 2, and also precalculus/trigonometry, and also physics. And also AP physics!

OpenStudy (timaashorty):

That's an amazing, wish you luck with you're teaching.(:

OpenStudy (timaashorty):

*your

OpenStudy (agent0smith):

Thanks! :D my students seem to like me!